Observers: Fuyon Bian and Peter Milne

Operator: Aaron Ceranski

Summary

Did not open due to problem with building rotation.

Details

All Times UT

Fifth Night of the AZ queue. (last night)
0221 Telescope motion problem, can only move in elevation.
0530 Work continues on this problem, no observations yet.
0635 Ditto.....
0800 Problem cannot be fixed tonight, we are closed and done.

-- XiaohuiFan - 31 Mar 2009
Topic revision: r5 - 31 Mar 2009, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback